PREDISSOCIATION DYNAMICS OF THE T-SHAPED AND LINEAR ISOMERS OF I_{2}(B)Ar

\underline{AMY BURROUGHS} and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Three-dimensional quantum dynamics models for the predissociation of I_{2}(B)Ar were examined by Gray and Roncero.^{a} Their calculations predicted highly structured rotational population distributions for the I_{2}(B) products. The distributions were characteristic of dynamics controlled by intramolecular vibrational energy redistribution (IVR). To test this model we have examined final state distributions resulting from predissociation of T-shaped I_{2}(B)Ar using optical-optical double resonance (OODR) techniques. We observed smooth rotational distributions that are indicative of direct dissociation or IVR occurring with level densities near the statistical limit.

OODR measurements were used to probe the highest vibrational level of I_{2}(B) populated by predissociation of Tshaped I_{2}(B)Ar. This is a matter of interest as Stevens Miller et al.^{b} have suggested that previous measurements^{c} may be in error, leading to overestimation of the bond strength for the B and X states of I_{2}Ar. The previous measurements relied on detection of dispersed fluorescence from I_{2}(B), and may have been influenced by non-adiabatic decay channels. OODR is a more sensitive technique, potentially capable of detecting the highest energy products prior to non-radiative decay. However, the OODR experiments yield results that are in agreement with the earlier study.c The data do not resolve conflicts associated with the relative bond strengths of the T-shaped and linear isomers.c

OODR techniques are being used to characterize the predissociation of linear I_{2}(B)Ar. Initial measurements show that I_{2}(B) from the linear isomer has less rotational energy than I_{2}(B) from the T-shaped complex.

Work supported by the National Science Foundation

aS. K. Gray and O. Roncero J. Phys. Chem. \underline{99}, 2512 (1995)

bA. E. Stevens Miller, C. Chuang, H. C. Fu, K. Higgins and W. Klemperer J. Chem. Phys. \underline{111}, 7844 (1999)

cJ. A. Balzy, B. M. DeKoven, T. D. Russell and D. H. Levy J. Chem. Phys. \underline{72}, 2439 (1980)